5 years ago sean was 3 times as old as his brother. today he is 2 times as old as his brother.how many years older than his brother is sean? how old is sean now?
Answer by scott8148(6628)

"5 years ago sean was 3 times as old as his brother" ___ s - 5 = 3(b - 5) ___ s - 5 = 3b - 15

"today he is 2 times as old as his brother" ___ s = 2b

2b - 5 = 3b - 15 ___ 10 = b

s = 2b = 20